### v2

# Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty

Abstract
DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline. In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image. This raises the question: can we trust and use all of these predictions? Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions. More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated. Based on these insights, we demonstrate that identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
Building on this viewpoint, we first address the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error. Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized. In response, we present Object-level Calibration Error (OCE), which assesses the calibration quality more effectively and is suitable for both ranking different models and identifying the most reliable predictions within a specific model. As a final contribution, we introduce a post hoc uncertainty quantification (UQ) framework that predicts the accuracy of the model on a per-image basis. By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, our framework assesses the reliability of the DETR model for each test image.
(a) Random boxes. Low Confidence.
(b) Accurate boxes. Calibrated.
(c) Accurate boxes. Uncalibrated.
Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated. This figure illustrates how DETR can handle the remaining predictions (i.e., red and gray boxes). In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with accurate prediction for the bounding boxes, or (c) poorly calibrated confidence with accurate boxes. Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
I Introduction
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [1, 2, 3, 4, 5, 6] to identify and locate objects within images. More recently, DEtection TRansformer (DETR) [7] has revolutionized the field by utilizing a Transformer encoder-decoder architecture to offer a scalable end-to-end prediction pipeline, where the model predicts a set of bounding boxes and class probabilities. This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks. While notable progress has been made, the reliability of these predictions remains insufficiently investigated.
The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used. For example, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions. As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler. Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image. However, owing to DETR’s unique set-prediction mechanism, quantifying its image-level reliability is not straightforward and remains largely underexplored.
Moreover, there is an ongoing debate within the community about whether DETR can be considered an entirely end-to-end solution, free from any post-processing requirements in practice. In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used. Practitioners often employ heuristic approaches to determine this subset, such as by setting a user-defined threshold to retain only a small subset of high-confidence (e.g., $>0.7$) outputs, as seen in the official demo. Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence score. Nonetheless, how different subset selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remain under-explored, and their significance has yet to be fully recognized.
Our primary findings reveal that DETR predictions within the same image are interdependent, leading to significantly varying levels of reliability. Since DETR is trained using gradients derived exclusively from optimally matched predictions (i.e., the predictions most closely corresponding to each annotated object), it will output at least one well-calibrated prediction for each perceived object while having flexibility in handling the remaining unmatched predictions. Throughout the paper, we refer to those matched predictions as positive predictions and the others as negative predictions. In principle, to avoid an inaccurate match, DETR could assign those unmatched predictions low confidence with random bounding boxes (Figure 1a). Alternatively, it could generate similarly accurate bounding boxes with (nearly) equally calibrated confidence (Figure 1b). Otherwise, it may assign poorly calibrated confidence scores (Figure 1c). Our analysis indicates that a prediction generated by DETR mostly follows the third scheme.
Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and addresses the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability? (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set? and (RQ3) How can we accurately assess DETR’s image-level reliability? In addressing these questions, our specific contributions include:
  1. 1.
    We provide both qualitative and quantitative insights into how positive and negative predictions influence the model’s reliability. Our analysis reveals that, unlike positives, negative predictions are often poorly calibrated and therefore should be handled separately. Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use (Section IV).
  2. 2.
    Identifying ground-truth positives and negatives ideally requires a matching process based on ground-truth annotations, which are unavailable at test time. Consequently, an alternative framework for identifying positive predictions is needed. However, we show that existing performance and calibration metrics—such as average precision [16, 17, 18] and expected calibration error [8, 19, 20]—are inadequate for this purpose. To address this, we introduce a new calibration metric, object-level calibration error (OCE), which assesses calibration error along ground-truth objects rather than predictions. We demonstrate that OCE is more suitable for both ranking different models by their calibration qualities and identifying the positive predictions (Section V).
  3. 3.
    Based on our findings, we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric. We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios (Section VI).
In addition to addressing the three primary research questions, we present a comparative analysis of standard post-processing methods for identifying positive predictions in DETR from an uncertainty quantification (UQ) perspective (Section VII). Our results indicate that thresholding on the predictive confidence score is more effective than widely used techniques such as top-$k$ selection or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
II Related Work
Calibration. Model calibration refers to how well a model’s predicted confidence scores align with the actual likelihood of those predictions being correct. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. For instance, if a model predicts an event with 80% confidence, it should be correct around 80% of the time for that event to be considered well-calibrated. Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
To evaluate the alignment, they measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction. To address these challenges, D-ECE [8] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$). On the other hand, LaECE [21] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
However, as we demonstrate, DETR exhibits poor calibration on negative predictions, regardless of its actual reliability. In practice, users often focus only on positive samples. Yet, when model calibration is evaluated across the entire set of predictions, including the negatives, the resulting evaluation becomes highly biased and therefore unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in later sections, a threshold of $0.3$ approximates the optimal positive predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable. However, the optimal threshold is not guaranteed to be always $0.3$ for other models like UP-DETR, which introduces potential risks. In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Uncertainty Quantification. Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, [22] propose a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. [23] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, [19] employ an auxiliary detection model capable of expressing its confidence. Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing UQ techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models. In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs. Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
III Preliminaries
Figure 2: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are sent to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
III-A Detection Transformer (DETR)
We consider a test image $\bm{x}$ and denote $\mathcal{D}_{\bm{x}}$ as the set of ground truth objects present in the image. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{D}}_{\theta}(\bm{x})$. Each prediction $d\in\hat{\mathcal{D}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. We refer to Figure 2 for an illustration.
It is noteworthy that the encoder follows the common structure of standard computer vision models, whose reliability has been relatively widely explored [26, 27, 28, 29]. This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [30]. However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction. Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of DETR.
III-B Bipartite Matching
Since the number of queries in DETR, $M$, is much higher than the number of annotated objects, DETR matches each object with the corresponding best model prediction during its training. To compute this optimal matching for the predictions in a given image, a bipartite matching algorithm is applied. More specifically, a matching cost between each pair of a given prediction and an object is defined as follows:
$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (1)
where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$. $\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [31] loss between bounding boxes. After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm [32, 7].
III-C Expected Calibration Error (ECE) Metrics
D-ECE. For a given object detection model, detection expected calibration error (D-ECE) [8] quantifies how closely the model’s predicted confidences align with its observed precision. Specifically, let $\hat{\mathcal{D}}$ be the set of all detections that the model produces on the validation set. Each detection has an associated predicted confidence score $\hat{p}\in[0,1]$. To compute D-ECE, the confidence space $[0,1]$ is partitioned into $J$ bins, for instance via uniform intervals $\big{[}0,\tfrac{1}{J}\big{)},\big{[}\tfrac{1}{J},\tfrac{2}{J}\big{)},\dots,\big{[}\tfrac{J-1}{J},1\big{]}$. Let $\widehat{\mathcal{D}}_{j}\subseteq\widehat{\mathcal{D}}$ be the subset of detections falling into the $j$-th bin. Denote by $\bar{p}_{j}$ the average confidence of detections in bin $j$, and by $\mathrm{precision}(j)$ the empirical precision in bin $j$. A detection is considered a true positive if it has the correct predicted class label and its IoU with the ground-truth box exceeds a given threshold $\tau$. The empirical precision $\mathrm{precision}(j)$ is thus the fraction of detections in $\widehat{\mathcal{D}}_{j}$ that are true positives.
D-ECE then aggregates the absolute difference between each bin’s average confidence and observed precision, weighted by bin size, and is defined as:
[…]
Following the standard protocol [20], two IoU thresholds ($\tau=0.5$ and $\tau=0.75$) are used, and the D-ECE scores obtained under these thresholds are averaged to yield the final metric. Because this process is carried out on a validation set, it provides an overall measure of the model’s calibration performance on that set. A lower D-ECE indicates that the confidence scores are better aligned with actual precision, while a higher D-ECE reveals a potential mismatch between confidence and true detection performance.
LaECE. To account for both localization error and classification precision, localisation-aware expected calibration error (LaECE) [19] aligns model confidence with the product of the predicted bounding box’s precision and its IoU with the ground-truth box. Specifically, LaECE is computed in a class-wise manner to mitigate class imbalance. Let $c$ index the class, and let $\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$. Partition the confidence scores into $J$ bins, and let $\widehat{\mathcal{D}}_{j}^{c}\subseteq\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$ in bin $j$. Define $\bar{p}^{c}_{j}$ to be the average confidence in $\widehat{\mathcal{D}}_{j}^{c}$, $\mathrm{precision}^{c}(j)$ to be the corresponding precision, and $\bar{\mathrm{IoU}}^{c}(j)$ to be the average IoU of $\widehat{\mathcal{D}}_{j}^{c}$. The LaECE metric is then given by:
$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)\Bigr{|}$ (3)
where $K$ is the total number of classes, and an IoU threshold $\tau$ is employed to determine whether a detection is considered a true positive.
LaECE0 [20] sets the IoU threshold $\tau$ to zero and assigns an average IoU of zero to false positive detections, thereby reducing the original formulation. In particular, if a detection is not a true positive under $\tau=0$, its IoU contribution is taken to be zero. The resulting metric becomes:
$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)\Bigr{|}.$ (4)
By setting $\tau=0$, false positives have $\bar{\mathrm{IoU}}=0$, and thus the calibration criterion focuses more explicitly on the contribution of bounding box localization to the overall calibration error.
III-D Image-Level Reliability
We introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image, extending [33, 29].
Definition 1.
We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{D}}_{\theta}(\bm{x}),~{}\mathcal{D}_{\bm{x}}\big{)}$ (5)
where one can use any standard performance metrics, such as average precision and recall, for $\mathsf{Perf}$ depending on the user’s requirements.
By its definition, image-level reliability directly addresses the model’s applicability to a given test instance. However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible. Therefore, this paper’s objective is to develop a method that assigns a quantitative score to each image instance, closely aligning with the model’s image-level reliability.
(a) Predictions within High-Reliability Image
(b) Predictions within Low-Reliability Image
Figure 3: Visualizations of the predictions generated by Cal-DETR. The optimal positive prediction (indexed by 0 and bordered in blue) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented. For each prediction and layer, the maximum confidence score and its corresponding label are visualized. When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low despite its accurate bounding box predictions. On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration. However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
IV Impact of Negative Predictions on DETR’s Reliability
IV-A Motivation and Scope
Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{D}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects. To address this issue during model training, a bipartite matching algorithm is used to find the optimal matching prediction [7] for each ground truth object based on the alignment of the class label and bounding box (as detailed in Section III-B). The training loss depends on the quality of this matching, and as a result, the parameters $\theta$ are primarily optimized to enhance the accuracy of these matched predictions. In this paper, we refer to the matched predictions as optimal positive predictions, while the remaining predictions are termed optimal negative predictions.
At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive predictions remain unknown. This raises the question of whether all predictions are trustworthy, or should only a specific subset be chosen—and if so, which subset? If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 1b), the large number of predictions would not be a concern. We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive prediction per object. Meanwhile, unmatched negative predictions, which often have accurate bounding boxes, receive uncalibrated low scores (e.g., Figure 1c). As a result, the overall calibration error becomes significantly high if we include those noisy predictions in the final inference results. Hence, as detailed in the following sections, improper handling of negative predictions can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios. For this reason, properly distinguishing the positives from the negatives is a crucial task for ensuring the reliable use of DETR.
IV-B Exploring the Anatomy of DETR’s Predictions
We begin our analysis by visualizing and examining the outputs generated from the DETR decoders. Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward. We address this by reapplying the final feedforward network that operates on the last layer, $f_{\phi}$, to the intermediate layers. This allows us to transform each representation vector into its associated bounding box and class label. This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [34]. Sample visualizations are in Figure 3.
In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions. In this early stage, the distinction between positive and negative queries can be ambiguous. However, the self-attentions through the subsequent decoder layers progressively refine these predictions. By the final layer, the model selects a single query (i.e., optimal positive prediction) and assigns a confidence score based on its understanding of the image and the object. In contrast, the confidence scores for neighboring queries (i.e., optimal negative predictions) do not increase to the same extent as the positives and even decrease in high-reliability images. In contrast, in low-reliability images, the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged. Based on this observation, we present our claim:
Claim 1.
Predictions from DETR within a given image exhibit varying levels of reliability. For each object in the image, the optimal positive prediction is calibrated (i.e., reliable), while the remaining optimal negative predictions are not.
IV-C Numerical Analysis
(a) Object-level Calibration Error
(b) Correlation to Image-level Reliability
Figure 4: A visualization of the difference in calibration between positive and negative predictions on the COCO dataset. In Figure 4a, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions. In Figure 4b, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
Setup. To provide quantitative support for our claim regarding the varying levels of reliability across predictions, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR), Cal-DETR [12], and DINO. Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set. The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). Further details can be found in the Appendix.
Object-level Calibration Error. We evaluate and compare the object-level calibration errors on the optimal positive predictions and the optimal negative predictions. The optimal positives and negatives are determined using the previously mentioned bipartite matching algorithm.
As illustrated in Figure 3, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy. Furthermore, as shown in Figure 4a, the optimal positive predictions are significantly better calibrated than the negative ones. This trend is evident across different state-of-the-art DETR variants and datasets, all of which exhibit poor calibration quality for optimal negative predictions.
Correlation to Image-Level Reliability. We compute $\mathsf{ImReli}$ for each image by using the same COCO evaluator [18] to obtain the image-wise AP score for $\mathsf{Perf}$ in Equation 1, passing the predictions and annotations for each image individually rather than the entire image set. Then, we measure the Pearson correlation coefficient (PCC) between the $\mathsf{ImReli}$ and the average confidence score of the optimal positives and the optimal negatives, respectively.
As shown in Figure 4b, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models. Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability. This finding reinforces our claim, highlighting the importance of distinguishing between positive and negative predictions, especially when conducting image-level UQ in DETR. Further visualizations for the other datasets are provided in the Appendix.
V A Systematic Framework for Identifying Positive Predictions
As illustrated so far, using positive predictions is crucial for the reliable use of DETR. Nonetheless, due to the lack of annotations, it is not feasible to identify either the ground truth or the optimal positive predictions during the inference stage. Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability. Given that the optimal matching achieves a small calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
Figure 5: Impact of confidence threshold selection on various performance and error metrics in Cal-DETR evaluated on COCO. Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
V-A Limitations of Existing Metrics
To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and localization recall precision (LRP) [21, 36], for identifying positive predictions. Specifically, we sweep the threshold for the confidence score to generate different subsets of DETR predictions (i.e., the predicted positives). We then assess the performance of each metric for these subsets and determine the threshold value that yields the highest performance. Analytical results using Cal-DETR are illustrated in Figure 5. For definitions of AP and LRP and additional results on the other DETR models, please refer to the Appendix.
Primarily, as noted in several studies [21, 19, 20, 36] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$ because AP does not penalize harshly for having low confidence predictions. However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications. Furthermore, using hundreds of predictions diminishes the interpretability of the model.
In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. This is a structural pitfall that prediction-level ECEs commonly face [20]. Since ECEs do not penalize missed detections (i.e., false negatives), they could achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions. Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when paired with a large confidence threshold.
On the other hand, LRP, a localization-focused performance metric, could be used instead. However, LRP is not designed as a calibration metric so does not explicitly consider calibration error. As a result, the model ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities, as shown in the following numerical analysis.
V-B Proposed Metric: Object-Level Calibration Error
Notation. For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider a set of ground truth object annotations $\mathcal{D}_{i}=\{d_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}}$, and a set of DETR predictions: $\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})=\{\hat{d}_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries. Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
Definition 2.
Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{A}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{D}}_{\theta}$. We define an object-level calibration error (OCE) as the average Brier score per object:
$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}$ (6)
$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\bar{p}_{i,j}[c]\big{)}^{2}$ (7)
$\bar{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (8)
where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $d_{i,j}$ and we propose two variants:
$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\delta\}~{}$ $(\mathsf{OCE_{ENS,\delta}})$ (9)
$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmax}_{q}\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (10)
The difference is that $\mathsf{OCE_{ENS,\delta}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching. When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$. Following [18, 20], we use IoU thresholds of $\delta=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
The introduced calibration error has two desirable characteristics. First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects. Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects. This is achievable due to the primary difference between our OCE and ECEs: while ECEs are evaluated based on selected predictions, OCE is assessed along the ground truth objects. This ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics. Thus, this metric can effectively assess not only whether the given subsets $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compared to existing metrics.
V-C Numerical Analysis: Effectiveness of OCE
Table 1: Comparison of the effectiveness of the proposed OCE and exisiting metrics, focusing on their correlation with the models’ overall calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, different post-processing schemes (i.e., methods for identifying positive predictions) are applied and specified in parentheses. These include using the top 100 predictions or choosing predictions with confidence scores above the threshold that optimizes the respective metric. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE consistently provides strong correlations with optimal calibration errors and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
[…]
References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980 $\pm$ 0.039
[…]
$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970 $\pm$ 0.050
Baselines AP (Top-100) [17, 18] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634 $\pm$ 0.149
AP@50 (Top-100) [18] 0.380 0.409 0.164 0.687 0.710 0.606 -0.360 0.079 -0.440 0.249 $\pm$ 0.402
AP@75 (Top-100) [18] 0.652 0.680 0.458 0.830 0.829 0.757 0.765 0.930 0.705 0.734 $\pm$ 0.127
LRP (LRP) [21] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137 $\pm$ 0.257
D-ECE (D-ECE) -0.757 -0.700 -0.772 0.365 0.284 0.337 0.309 0.436 0.336 -0.018 $\pm$ 0.514
D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144 $\pm$ 0.614
D-ECE (0.3) [8] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928 $\pm$ 0.067
D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895 $\pm$ 0.116
D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914 $\pm$ 0.095
LA-ECE0 (LA-ECE0) -0.960 -0.930 -0.968 -0.913 -0.874 -0.907 -0.826 -0.839 -0.835 -0.895 $\pm$ 0.051
LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697 $\pm$ 0.228
LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639 $\pm$ 0.254
LA-ECE0 (LRP) [20] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854 $\pm$ 0.097
LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810 $\pm$ 0.112
Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951 $\pm$ 0.078
$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934 $\pm$ 0.086
To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration errors across four DETR models. As baselines, we measure ECEs using the top $100$ detections  a standard confidence threshold of 0.3, and the threshold that minimizes the different metrics on the validation set. In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set. For a fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set. We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the PCC correlation to the other scoring schemes across different datasets.
Table 1 presents the correlations between the three calibration metrics on the optimal positive set and various methods. The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—are highly correlated. This indicates that all three metrics effectively capture the notion of calibration quality. Additionally, these metrics show a decent correlation with AP metrics. It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
LRP, D-ECE, and LA-ECE0 exhibit inverse correlations with calibration quality when used alone. Furthermore, LA-ECE0 with a non-optimal positive set also shows an inverse correlation with the overall calibration errors estimated on the optimal positive set. This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
Another important attribute of OCE is that it can not only compare different models but also identify the most reliable configuration of the same model. As illustrated in Figure 5, the confidence threshold vs. OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$. This is particularly notable because it reflects the practical choices in many calibration studies to date [12, 20], affirming the reliability of using OCE.
The D-ECE and OCE scores measured at the optimal LRP threshold—denoted by D-ECE (LRP) and OCE (LRP), respectively—show a strong correlation to those computed from the optimal positive set; thus, the method of using LRP to identify a positive set [20] appears acceptable. Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
Nonetheless, the biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed. While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference. Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
Using a fixed separation threshold of $0.3$ might serve as a reasonable approximation for distinguishing between optimal positives and negatives. However, trained or post-calibrated DETR models often have varying optimal confidence thresholds. For example, UP-DETR has an optimal threshold of approximately $0.5$, and low-temperature calibrated models are likely to require even higher thresholds. Consequently, employing a fixed threshold poses potential risks. Similarly, using a fixed number of top confident samples (e.g., Top-$100$) is inadequate, as it may include more predictions than the actual number of ground truth positives.
VI Quantifying Image-level Reliability
As demonstrated, positive and negative predictions exhibit varying levels of reliability. Interestingly, having predictions with low confidence scores does not necessarily imply low reliability. Our empirical observations show that confidence scores in negative predictions are actually inversely correlated with image-level reliability. More specifically, for a reliable instance, we observe that the confidence of positive predictions increases across the decoder layers, while that of negative predictions remains low; this results in a large gap between positive and negative predictions. In contrast, for unreliable instances (e.g., Figure 3a), the confidence score of the positive prediction does not increase across the layers, whereas the scores of negative predictions are either slightly elevated or remain unchanged. Consequently, there is a small gap between positive and negative predictions (e.g., Figure 3b).
VI-A Proposed Method: Quantifying Reliability by Contrasting
Based on the finding, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
$\mathsf{ContrastiveConf}(\bm{x})=\mathsf{Conf}^{+}(\bm{x})-\mathsf{Conf}^{-}(\bm{x})$ (11)
$\mathsf{Conf}^{+}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (12)
$\mathsf{Conf}^{-}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (13)
where $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor that can be determined on the validation set. We include an ablation study of the scaling factor in the subsequent section. Our results demonstrate that the proposed method is robust to the choice of scaling factor and consistently outperforms baseline approaches, achieving the robust performance with a scaling factor of $5.0$ $10.0$.
Identifying a $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{D}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach. In practice, however, neither the ground truth nor the optimal positives are available during the test time. Instead, we approximate the ground truth separation by applying a post-processing algorithm $\mathcal{A}^{*}$ that minimizes the calibration error on the validation set:
$\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})=\mathcal{A}^{*}\circ\hat{\mathcal{D}}_{\theta}(\bm{x})$ (14)
$\mathcal{A}^{*}=\text{argmin}_{\mathcal{A}}\mathsf{OCE}(\mathcal{A}\circ\hat{\mathcal{D}}_{\theta};\mathcal{I}_{val})$ (15)
where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{N_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$.
Table 2: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (5), with different models. $\mathsf{Perf}$ is evaluated using AP. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), or the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed contrasting method, $\mathsf{ContrastiveConf}$, achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.503 0.618 0.670 0.635 0.555 0.647 0.649 0.633 0.561 0.642 0.662 0.634
$\mathsf{Conf}^{-}$ (Optimal) -0.608 -0.584 -0.572 -0.586 -0.293 -0.296 -0.330 -0.311 -0.177 -0.235 -0.212 -0.196
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.648 0.684 0.662 0.601 0.659 0.656 0.644 0.612 0.660 0.667 0.647
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
$\mathsf{Conf}^{+}$ (0.3) 0.476 0.464 0.539 0.504 0.229 0.353 0.385 0.399 0.195 0.256 0.258 0.258
Ours $\mathsf{Conf}^{+}$ (OCE) 0.477 0.478 0.547 0.512 0.252 0.330 0.393 0.370 0.191 0.276 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
VI-B Numerical Analysis
We compare different methods for quantifying per-image reliability based on different separation methods. For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding. We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed based on AP metric (see Appendix for the detail).
Table 2 shows that the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability. Interestingly, the absolute value of $\mathsf{Conf}^{-}$ often surpasses that of $\mathsf{Conf}^{+}$ when a non-optimal separation scheme is applied. This observation highlights how our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$, achieves robust performance across diverse settings.
(a) COCO
(b) Cityscapes
(c) Foggy Cityscapes
Figure 6: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE). Pearson correlation coefficient (PCC) using various scaling factors is reported. The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
In addition, we conduct an ablation study on the scaling factor ($\lambda$), with results presented in Figure 6. The study reveals that the best performing scaling factor for COCO dataset lies between $5.0$ and $10.0$. Notably, this range remains effective even for out-of-distribution datasets such as Cityscapes and Foggy Cityscapes. Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line). However, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes. Thus, caution is needed when applying to substantially different datasets.
Table 3: Comparison of post-processing schemes based on calibration quality (i.e., OCE). We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. The optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Thresholding 0.276 (0.35) 0.450 (0.20) 0.416 (0.25) 0.436 (0.25) 0.313 (0.45) 0.430 (0.30) 0.413 (0.30) 0.402 (0.35) 0.387 (0.40) 0.488 (0.30) 0.469 (0.30) 0.458 (0.30)
Top-$k$ 0.358 (20.00) 0.485 (20.00) 0.457 (20.00) 0.479 (20.00) 0.352 (20.00) 0.451 (20.00) 0.436 (20.00) 0.422 (20.00) 0.416 (20.00) 0.500 (20.00) 0.490 (20.00) 0.473 (20.00)
NMS 0.358 (0.90) 0.535 (0.90) 0.624 (0.90) 0.521 (0.90) 0.426 (0.90) 0.625 (0.90) 0.784 (0.90) 0.583 (0.90) 0.495 (0.90) 0.650 (0.90) 0.794 (0.90) 0.609 (0.90)
Table 4: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}(\lambda=1.0)$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds often exceed $0.5$, highlighting the potential limitation of using a fixed threshold for image-level UQ.
UQ Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
$\mathsf{Conf}^{+}$ Thresholding 0.479 (0.25) 0.478 (0.20) 0.547 (0.25) 0.512 (0.25) 0.333 (0.80) 0.355 (0.25) 0.411 (0.20) 0.403 (0.35) 0.229 (0.50) 0.291 (0.25) 0.281 (0.15) 0.282 (0.35)
Top-$k$ 0.067 (1.00) 0.132 (1.00) 0.252 (1.00) 0.155 (1.00) -0.037 (2.00) 0.019 (1.00) 0.078 (1.00) 0.021 (1.00) 0.080 (1.00) -0.006 (1.00) 0.026 (1.00) 0.090 (1.00)
NMS -0.524 (0.10) -0.508 (0.10) -0.488 (0.10) -0.510 (0.10) -0.370 (0.10) -0.324 (0.10) -0.273 (0.10) -0.338 (0.10) -0.205 (0.10) -0.251 (0.10) -0.202 (0.15) -0.204 (0.10)
$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.513 (0.25) 0.563 (0.25) 0.540 (0.25) 0.408 (0.80) 0.389 (0.30) 0.422 (0.20) 0.437 (0.35) 0.312 (0.75) 0.317 (0.25) 0.288 (0.15) 0.311 (0.35)
Top-$k$ 0.629 (1.00) 0.268 (1.00) 0.320 (1.00) 0.290 (1.00) 0.409 (1.00) 0.169 (1.00) 0.151 (1.00) 0.220 (1.00) 0.284 (1.00) 0.077 (1.00) 0.060 (1.00) 0.174 (1.00)
NMS -0.227 (0.10) -0.157 (0.90) -0.402 (0.10) -0.140 (0.90) -0.044 (0.80) -0.034 (0.75) -0.203 (0.10) -0.080 (0.80) 0.035 (0.75) -0.035 (0.75) -0.149 (0.90) -0.075 (0.70)
(a) Thresholding
(b) Top-$k$
(c) NMS
Figure 7: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
VII Comparative Study on Post-Processing
In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality and UQ performance. For the post-processing (i.e., separation) methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the Top-$k$ predictions, and (3) utilizing NMS.
Primarily, we evaluate OCE by varying the hyperparameter for each method (i.e., confidence threshold, $k$, and IoU threshold, respectively) and compare the resulting performance. In this first analysis, to exclude the dependency of the final performance on OCE, we choose the best hyperparameter on the validation set for each setting and compare their best possible performances. Table 3 shows that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. Top-$k$ becomes optimal when $k=20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 4. In this experiment, we fixed the scaling factor $\lambda$ at $1.0$ to minimize its influence on selecting the optimal hyperparameter. We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications. Moreover, the correlation is often negative when NMS is applied along with the $\mathsf{Conf}^{+}$ framework. This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times. As a result, applying NMS leads to an inaccurate reliability assessment. However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework; even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative predictions.
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. Figure 7 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also present this with exemplary visualizations in Figure 8 and in the Appendix.
(a) Thresholding (0.1)
(b) Thresholding (0.3)
(c) Thresholding (0.9)
(d) Top-$1$
(e) Top-$20$
(f) Top-$100$
(g) NMS ($\tau=0.1$)
(h) NMS ($\tau=0.5$)
(i) NMS ($\tau=0.9$)
Figure 8: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes. Optimal positive and negative predictions are highlighted with green and red boxes, respectively. As shown, the top-$k$ and NMS approaches often include too many negative predictions, degrading the calibration quality. Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions. Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.
VIII Conclusion
The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks. In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated positive predictions. To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these positive predictions effectively. Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability. We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
Acknowledgments
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
Appendix A Dataset
For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments. Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set. Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO. More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly. In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO. The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g. building, sky, sidewalk). The pre-trained model weights were obtained from their respective official implementations.
Appendix B Performance Metrics
Average precision (AP) [17] is the most common accuracy metric for object detectors. Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded. Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough. The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
where $p(r)$ is calculated by the maximum precision at recall level $r$.
[…]
$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object. LRP is then computed by averaging LRP0.5 and LRP0.75.
Appendix C Omitted Experimental Results
This section provides the omitted experimental results. Figure 10 extends Figure 5 in Section V-A across different DETR variants and datasets. Figure 11 extends Figure 8 in Section VII across different DETR variants and post-processing schemes.
Appendix D Contrasting with Confounding Negatives for Image-level UQ
In Table 2, we observe an interesting empirical result: $\mathsf{Conf}^{+}$ using the top 100 predictions (i.e., the fourth row) exhibits a stronger negative correlation compared to $\mathsf{Conf}^{-}$ using the optimal negative predictions (i.e., the second row). While a stronger negative correlation might not be desirable when used alone for UQ, this phenomenon can be leveraged within our contrasting framework, similar to how negative predictions are utilized.
For instance, rather than contrasting the average confidence scores of positive predictions against those of negative (non-positive) predictions, we can instead contrast them against the top 100 predictions. However, since positive predictions are included within this top 100 set, an alternative and potentially clearer approach is to contrast the positive predictions specifically against the top 100 non-positive predictions.
Nevertheless, employing a fixed value of $k=100$ might introduce unintended consequences, particularly depending on the number of ground truth objects present in an image. We hypothesize that among non-positive predictions, those with bounding boxes significantly overlapping the positive predictions play a more meaningful role. Consequently, negative predictions that have random bounding boxes and very low confidence scores (as illustrated in Figure 1a) can be safely disregarded during the contrastive evaluation.
To formalize this intuition, we introduce the concept of confounding negatives:
Definition 3.
Confounding negative predictions are defined as non-positive predictions whose bounding boxes significantly overlap with positive predictions, specifically exceeding a threshold overlap $\delta$:
$\hat{\mathcal{D}}_{\theta}^{-\delta}(\bm{x})=\Big{\{}d^{-}\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})\mid\max_{d^{+}\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\mathrm{IoU}(d^{-},d^{+})\geq\delta\Big{\}}.$ (18)
We conduct an ablation study comparing the original approaches, which use the entire set of negative predictions, against variants that employ different subsets of negatives described above. Our findings are summarized in Table 5.
Overall, we see that restricting attention to negatives with significant overlap can sharpen the negative correlation for $\mathsf{Conf}^{-}$, especially on in-distribution data (COCO), supporting our hypothesis that misleading or “confouding” boxes are more informative for contrastive scoring. On the other hand, using entire negatives sometimes proves more robust in near-OOD and OOD scenarios, particularly for $\mathsf{ContrastiveConf}$.
Table 5: Comparison of correlation results under $\mathsf{Conf}^{-}$ (rows 1–5) and $\mathsf{ContrastiveConf}$ (rows 6–10) when using different negative sets on three datasets: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). “Entire Negatives” refers to all non-positive boxes, while “Top-100 Negatives” and “Top-100 Positives” limit the selection to the 100 highest-confidence predictions in each category. Confounding negatives restricts the negative set to boxes overlapping positive boxes by at least $\delta$. Stronger (more negative) correlation values under $\mathsf{Conf}^{-}$ appear when negatives more closely resemble the positive boxes, and $\mathsf{ContrastiveConf}$ tends to benefit from using such confounding negatives, especially in in-distribution scenarios.
Methods Negatives COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
$\mathsf{Conf}^{-}$ Entire Negatives -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
Top-100 Positives -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
Top-100 Negatives -0.640 -0.607 -0.607 -0.620 -0.349 -0.385 -0.415 -0.404 -0.256 -0.315 -0.272 -0.279
Confounding Negatives ($\delta=0.5$) -0.641 -0.640 -0.574 -0.647 -0.330 -0.351 -0.364 -0.386 -0.250 -0.260 -0.254 -0.261
Confounding Negatives ($\delta=0.75$) -0.635 -0.628 -0.561 -0.602 -0.385 -0.388 -0.384 -0.390 -0.285 -0.279 -0.269 -0.254
$\mathsf{ContrastiveConf}$ Entire Negatives 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
Top-100 Positives 0.632 0.610 0.604 0.619 0.412 0.379 0.388 0.404 0.270 0.277 0.257 0.247
Top-100 Negatives 0.656 0.612 0.629 0.633 0.359 0.401 0.439 0.431 0.275 0.332 0.299 0.308
Confounding Negatives ($\delta=0.5$) 0.653 0.641 0.623 0.654 0.340 0.371 0.434 0.418 0.268 0.284 0.314 0.296
Confounding Negatives ($\delta=0.75$) 0.647 0.630 0.613 0.619 0.388 0.406 0.448 0.418 0.298 0.301 0.326 0.286
(a) Object-level Calibration Error
(b) Object-level Calibration Error
(Foggy Cityscapes)
(c) Corr. to Image-level Reliability
(d) Corr. to Image-level Reliability
(Foggy Cityscapes)
Figure 9: A visualization of the difference in calibration between positive and negative predictions on the Cityscapes and Foggy Cityscapes datasets. In Figures 9a and 9b, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions. In Figures 9c and 9d, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
(a) UP-DETR (COCO)
(b) D-DETR (COCO)
(c) Cal-DETR (COCO)
(d) DINO (COCO)
(e) UP-DETR (City)
(f) D-DETR (City)
(g) Cal-DETR (City)
(h) DINO (City)
(i) UP-DETR (Foggy)
(j) D-DETR (Foggy)
(k) Cal-DETR (Foggy)
(l) DINO (Foggy)
Figure 10: Impact of confidence threshold selection on various performance metrics in UP-DETR, Deformable-DETR, Cal-DETR, and DINO on COCO, Cityscapes, and Foggy Cityscapes. Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
(a) Thresholding (UP-DETR)
(b) Top-$k$ (UP-DETR)
(c) NMS (UP-DETR)
(d) Thresholding (D-DETR)
(e) Top-$k$ (D-DETR)
(f) NMS (D-DETR)
(g) Thresholding (Cal-DETR)
(h) Top-$k$ (Cal-DETR)
(i) NMS (Cal-DETR)
(j) Thresholding (DINO)
(k) Top-$k$ (DINO)
(l) NMS (DINO)
Figure 11: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability across different DETR models on COCO for different post-processing schemes.